# Ontological Foundations for Contrastive Explanatory Narration of Robot Plans

### v1

@@ Abstract
[iriAffiliation]organization=Institut de Robòtica i Informàtica Industrial, CSIC-UPC,addressline=Llorens i Artigas 4-6, city=Barcelona, postcode=08028, country=Spain
@@
A set of logical rules (axioms) is proposed to automate the reasoning behind the comparison of plans and to validate the ontology
It is presented a novel algorithm to construct contrastive explanatory narratives exploiting the ontology and the rules
The proposal is evaluated using several planning domains against a baseline method
@@ 2 Related work
Concerning the modeling of domain knowledge for reasoning, a common approach is to use sound formalisms such as formal ontologies. The literature shows that multiple ontologies have been lately developed and even standardized for different robotic applications and domains [6, 7, 8, 9]. The literature also encompasses various works that have concentrated on ontologically modeling concepts related to plans and their properties. In this regard, Bermejo-Alonso [10] conducted a survey of that domain, reviewing existing task and planning vocabularies, taxonomies, and ontologies, while also discussing their potential integration. The surveyed works mostly defined general concepts regarding plans, and only in some cases plan properties were discussed (e.g. cost and constraints). In another work, Moulouel et al. [11], proposed an ontology to model contextual knowledge in partially observable environments, which is used for probabilistic commonsense reasoning and probabilistic planning. They defined concepts to model the relationship between the context in which robot actions and observations take place and other elements used during the planning (e.g. reward function). These works are relevant in the domain, but if one wants to contrast plans during robots’ decision making, there is still a need for an ontological theory that formalizes the properties of plans.
@@
Ontologies have already shown potential to become the integrative framework for explainable robots [17]. However, there are other formalisms that could complement ontologies enabling other types of reasoning. For example, one might use non-monotonic logic, which supports abductive reasoning and is devised to capture and represent defeasible inferences, i.e., a kind of inference in which reasoners draw tentative conclusions, enabling reasoners to retract their conclusion(s) based on further evidence. Sridharan [16] proposed to integrate non-monotonic logic and deep learning, exploiting their synergy across representation, reasoning, and learning layers to operationalize the principles of explainable agency. Another example of complementary formalism is causal modeling, which supports reasoning to determine the relationship between causes and their effects. Love et al. [18], developed a system based on causal models to enable a robot to autonomously elicit interactions in multi-person environments, while generating real-time counterfactual explanations for its elicitation decisions. Those works presented interesting approaches to robot reasoning and explanation generation. In future work, we might consider the integration of our ontology-based methodology for explainable robots with those approaches.
@@ 3.3 Formalization of the model in OWL 2 DL
For practical use, the proposed ontological theory was formalized in OWL 2 DL. Hence, the axioms listed in this paper were translated to DL, more specifically, to the SROIQ(D) fragment of DL. It was possible to formulate each of the axioms in the target formalism with the exception of the value of plans’ qualities. Note that quality is often used as a synonym for property but not in DUL, where qualities are particulars and properties are universals. In this regard, DUL considers that ‘qualities inhere in entities’ [23]. Every entity (including qualities) comes with its own exclusive qualities, which exist as long as the entity exists. DUL distinguishes between a quality (the cost of a plan) and its value or quale (a numerical data value). Hence, when saying that two plans have the same cost, their costs have the same quale, but still they are distinct qualities. This is convenient to model and answer CQ2, since OWL 2 DL cannot model relationships between two data values or quales (i.e. one cannot state that 5 is a better cost than 10). However, one can model the relation between the qualities (e.g. ‘cost A’ has better quality value than ‘cost B’). Let’s consider that a plan ‘p’ has a cost ‘c’ whose value is ‘10’. Hence, the knowledge would be modeled as:
@@ 5.4 ACXON - An algorithm for contrastive explanatory ontology-based narratives
ACXON is a novel theoretical contribution, an algorithm that retrieves knowledge about divergences between ontological entities (e.g. plans), for later construction of textual contrastive explanatory narratives. The algorithm takes as an input a time-indexed knowledge graph $\mathcal{G_{T}}$ (as described in Sec. 5.3), the ontological class (or classes) of the pair of instances to narrate, the temporal locality (time interval of interest), and the level of specificity. Our focus is on contrastive narratives about Plans, but ACXON is general enough to work with other OWL 2 DL ontologies and classes. For instance, it might contrastively narrate the capabilities of a pair of agents (e.g. one can move for longer periods), or how two drinks are different to each other (e.g. one is healthier, tastier, etc.). Based on the level of specificity, there are three types of contrastive narratives. In this article, specificity refers to the amount of detail to be used during the narrative construction, i.e., the number of knowledge tuples.
@@ Acknowledgments
This work was partially supported by Project PID2020-118649RB-I00 funded by MCIN/ AEI /10.13039/501100011033; Project PCI2020-120718-2 funded by MCIN/ AEI /10.13039/501100011033 and by the ”European Union NextGenerationEU/PRTR”; Project EP/V062506/1 funded by EPSRC (UK); and also by the European Union under the project ARISE (HORIZON-CL4-2023-DIGITAL-EMERGING-01-101135959). G. Canal was supported by the Royal Academy of Engineering and the Office of the Chief Science Adviser for National Security under the UK Intelligence Community Postdoctoral Research Fellowship programme. A. Olivares-Alarcos was supported by the European Social Fund and the Ministry of Business and Knowledge of Catalonia through the FI 2020 grant.

### v2

@@ Abstract
[iriAffiliation]organization=Institut de Robòtica i Informàtica Industrial, Consejo Superior de Investigaciones Científicas (CSIC) - Universitat Politècnica de Catalunya (UPC),addressline=Llorens i Artigas 4-6, city=Barcelona, postcode=08028, country=Spain
@@
A set of logical rules (ontological axioms) is proposed to automate the reasoning behind the comparison of plans
A novel algorithm is presented to construct contrastive explanatory narratives exploiting the ontology’s structure
The proposal is evaluated against a baseline method using several planning domains
@@ 2 Related work
Concerning the modeling of domain knowledge for reasoning, a common approach is to use sound formalisms such as formal ontologies. The literature shows that multiple ontologies have recently been developed [25] for different robotic applications and domains. The literature also encompasses various works that have concentrated on ontologically modeling concepts related to plans and their properties. In this regard, Bermejo-Alonso [3] conducted a survey of that domain, reviewing existing task and planning vocabularies, taxonomies, and ontologies, while also discussing their potential integration. The surveyed works mostly defined general concepts regarding plans, and only in some cases plan properties were discussed (e.g. cost and constraints). In another work, Moulouel et al. [23], proposed an ontology to model contextual knowledge in partially observable environments, which is used for probabilistic commonsense reasoning and probabilistic planning. They defined concepts to model the relationship between the context in which robot actions and observations take place and other elements used during the planning (e.g. reward function). These works are relevant in the domain, but if one wants to contrast plans during robots’ decision making, there is still a need for an ontological theory that formalizes the properties of plans.
@@
Ontologies have already shown potential to become the integrative framework for explainable robots [24]. However, there are other formalisms that could complement ontologies enabling other types of reasoning. For example, non-monotonic logic supports abductive reasoning and is specifically designed to handle defeasible inferences. These allow reasoners to form tentative conclusions that can be retracted as new evidence emerges. Sridharan [32] proposed to integrate non-monotonic logic and deep learning, exploiting their synergy across representation, reasoning, and learning layers to operationalize the principles of explainable agency. Another example of complementary formalism is causal modeling, which supports reasoning to determine the relationship between causes and their effects. Love et al. [20], developed a system based on causal models to enable a robot to autonomously elicit interactions in multi-person environments, while generating real-time counterfactual explanations for its elicitation decisions. Those works presented interesting approaches to robot reasoning and explanation generation. In future work, we might consider the integration of our ontology-based methodology for explainable robots with those approaches.
@@ 3.3 Formalization of the model in OWL 2 DL
For practical use, the proposed ontological theory was formalized in OWL 2 DL. Hence, the axioms presented in this paper were translated to DL, in particular, to the SROIQ(D) fragment. Each of the axioms was translated in the target formalism with the exception of the value of plans’ qualities. Note that quality is often used as a synonym for property but not in DUL, where ‘qualities are particulars and properties are universals’. In this regard, DUL considers that ‘qualities inhere in entities’ [14]. Every entity (including qualities) comes with its own exclusive qualities, which exist as long as the entity exists. DUL distinguishes between a quality (the cost of a plan) and its value or quale (a numerical data value). Hence, when saying that two plans have the same cost, their costs have the same quale, but still they are distinct qualities. This is convenient to model and answer CQ2, since OWL 2 DL cannot model relationships between two data values or quales (i.e. one cannot state that 5 is a better cost than 10). However, one can model the relation between the qualities (e.g. ‘cost A’ has better quality value than ‘cost B’). Let’s consider that a plan ‘p’ has a cost ‘c’ whose value is ‘10’. Hence, the knowledge would be modeled as:
@@ 5.4 ACXON - An algorithm for contrastive explanatory ontology-based narratives
ACXON is a novel theoretical contribution, an algorithm that retrieves knowledge about divergences between ontological entities (e.g. plans), for later construction of textual contrastive explanatory narratives (see Alg. 1). The algorithm takes as an input a time-indexed knowledge graph $\mathcal{G_{T}}$ (as described in Sec. 5.3), the ontological class (or classes) of the pair of instances to narrate, the temporal locality (time interval of interest), and the level of specificity. Our focus is on contrastive narratives about Plans, but ACXON is general enough to work with other OWL 2 DL ontologies and classes. For instance, it might contrastively narrate the capabilities of a pair of agents (e.g. one can move for longer periods), or how two drinks are different to each other (e.g. one is healthier, tastier, etc.). Based on the level of specificity, there are three types of contrastive narratives. In this article, specificity refers to the amount of detail to be used during the narrative construction, i.e., the number of knowledge tuples.
@@ Acknowledgments
This work was partially supported by Project PID2023-152259OB-I00 funded by MCIN/ AEI /10.13039/501100011033; Project PCI2020-120718-2 funded by MCIN/ AEI /10.13039/501100011033 and by the ”European Union NextGenerationEU/PRTR”; Project EP/V062506/1 funded by EPSRC (UK); and also by the European Union under the project ARISE (HORIZON-CL4-2023-DIGITAL-EMERGING-01-101135959). G. Canal was supported by the Royal Academy of Engineering and the Office of the Chief Science Adviser for National Security under the UK Intelligence Community Postdoctoral Research Fellowship programme. A. Olivares-Alarcos was supported by the European Social Fund and the Ministry of Business and Knowledge of Catalonia through the FI 2020 grant.